In a National Achievement Test, suppose three test

booklets are tested at random. Let D represent the

defective test booklet and let N represent the

non-defective test booklet. Let X be the random

variable representing the number of non-defective

test booklets


1
Expert's answer
2021-03-26T13:50:19-0400

There are "2^3=8" possible outcomes

"DDD, DDN, DND, DNN," "NDD, NDN, NND, NNN"


Let "X" be the random variable representing the number of non-defective

test booklets. Then


"\begin{matrix}
 X & f \\
 0 & 1 \\
 1 & 3 \\
 2 & 3 \\
 3 & 1 \\
\end{matrix}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 & 3 \\ 
 \hdashline
 p(x) & 1/8 & 3/8 & 3/8 & 1/8
\end{array}"